OpenStudy (anonymous):

Enter the expression 2 cos 2(θ)−1, where θ is the lowercase Greek letter theta

OpenStudy (anonymous):

@hero

OpenStudy (anonymous):

@Australopithecus

OpenStudy (australopithecus):

Enter it into what?

OpenStudy (anonymous):

I have to enter it into a answer box

OpenStudy (australopithecus):

that expression looks ambiguous, I would assume it is this? \[2\cos(2\theta - 1)\] Not really understanding the purpose of this question

OpenStudy (anonymous):

Its just to use explicit expressions

OpenStudy (anonymous):

@Australopithecus

OpenStudy (australopithecus):

can you post a picture of the question? not really sure what it is asking still

OpenStudy (anonymous):

Did you get it?

OpenStudy (australopithecus):

oh yeah what I wrote isn't right

OpenStudy (anonymous):

Okay. Can you explain to me what's the right answer?

OpenStudy (australopithecus):

so they just want you to show that you know how to write: \[2\cos^2(\theta) - 1\]

OpenStudy (australopithecus):

From what I understand is this just a tutorial on how to use their mathematics notation software?

OpenStudy (anonymous):

Yes

OpenStudy (australopithecus):

click the buttons above

OpenStudy (australopithecus):

the square ones

OpenStudy (australopithecus):

or try

OpenStudy (australopithecus):

writing, 2cos^2(theta) - 1 you want whats in the box to look identical to whats written next to it

OpenStudy (australopithecus):

those buttons above should let you insert theta and there should be one that allows you to input exponents just play around with it you will get it

OpenStudy (anonymous):

What's theta?

OpenStudy (anonymous):

Wait so insert the same thing?

OpenStudy (dls):

or simply write cos2theta